Home‎ > ‎Crew 134‎ > ‎


GreenHab Officer Musilova to Cap Comm

The green hab is operating normally

The fan was set to 60 degrees at 8AM
The fan is operating normally at 68 degrees at 5PM

The heater was turned down to 40 degrees F at 8AM
The heater was turned up to 75 degrees F at 5PM

The pump is operating normally.

The plants haven't been watered today, since I watered them last night and the soil still looks damp. 

I have attached a picture of the plants in the zen room.

All the seedlings were returned to the upper hab at 5PM. 

Greenhab readings for 5PM today are as follows:

                         Current             Min                 Max
Temperature      68 Degrees F   46 Degrees F  95 Degrees F   
Humidity           17%                  11%                 19%

Meter            0.93 KwH
Duration        23h 50 min
Overflow         1.5 Liters    Removed Today at 5PM - Checked at 8AM and at 5PM

Best wishes,
Officer Musilova